OpenStudy (anonymous):

quadractic equation: x^2/3+2x=-2 can anyone verify my answer of x=-2+-i square root 2 is correct

OpenStudy (anonymous):

you can verify that, armed with the definition i^2 = -1, you get (1/3)(2 + i)^2 + 2(2 + i) + 2, if this equals 0 you're correct.

OpenStudy (anonymous):

ok. can you work the problem

OpenStudy (anonymous):

\[\frac{x^2}{3}+2x=-2\], this your question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{x^2+3(2x)}{3}=-2\] \[x^2+6x=-6\] \[x^2+6x+6=0\] From this point, we can see your answers are wrong because the discrimant is positive, meaning the quadratic has real number solutions

OpenStudy (anonymous):

bummer

OpenStudy (anonymous):

I will rework it.

OpenStudy (anonymous):

alright i'll let you know if you get the right answer

OpenStudy (anonymous):

since when is 6^2 - 4(6) <0 ? it's 36 - 24 = 12. positive discriminant, so your complex answer 2 - i cannot be correct.

OpenStudy (anonymous):

my answer is \[x=-3 \pm 3\sqrt{2}\]

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

try 1 more time?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

looks like you messed up evaluating the sqrt(12)

OpenStudy (anonymous):

\[-3 \pm 2\sqrt{3?}\]

OpenStudy (anonymous):

nope, want me to write the equation?

OpenStudy (anonymous):

or \[-3 \pm \sqrt{3}\]

OpenStudy (anonymous):

that ones right

OpenStudy (anonymous):

okay the 2 does cancel. thanks this has been very helpful

OpenStudy (anonymous):

yw